OpenStudy (anonymous):

80=1/9x^2+11/3x solve for x

OpenStudy (anonymous):

Do you mean \[80=\frac{1}{9}x^2+\frac{11}{3}x\]?

OpenStudy (anonymous):

If so, multiply everything through by the highest denominator to eliminate your fractions. Here, multiply through by 9 to get\[720=x^2+33x\]

OpenStudy (anonymous):

This is quadratic, so you have\[x^2+33x-720=0\] which can be factored into \[(x-15)(x+48)=0\]

OpenStudy (anonymous):

For this to be true, either one, or both, of the factors have to be zero. Hence, your x-values are\[x=15,x=-48\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

np